OpenStudy (anonymous):

the total value of a collection of dimes and nickels is $2.70. if the total number of coins is 39, how many of each coin is there?

OpenStudy (anonymous):

hold on im trying to figure it out

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

it is a pain

OpenStudy (anonymous):

yeaah lol

OpenStudy (anonymous):

\[\large \sf d+n=39 \\\large \sf .1d+.05n=2.7\]

OpenStudy (anonymous):

soooo, there is half a nickel??

OpenStudy (anonymous):

15 dimes 24 nickels?

OpenStudy (anonymous):

thanks man

OpenStudy (anonymous):

Woah there, no direct answers. @.Gjallarhorn.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

lets see.. 15 dimes is 1.50 24 nickels 1.20 addem and you get 2.70

OpenStudy (anonymous):

i know im explaining

OpenStudy (anonymous):

And the method you used to figure it out?

OpenStudy (anonymous):

15 dimes and 24 nickels add up to 39 coins :)

OpenStudy (anonymous):

That isn't a method, my friend :)

OpenStudy (anonymous):

i just used guess and eliminate

OpenStudy (anonymous):

and i got the answer and i explained why that was the answer by adding the values

OpenStudy (anonymous):

And that's the problem. We're here to help the asker UNDERSTAND, which your method doesn't do. Your method doesn't work on anything but things like this. With such a small margin of answers, it's easy to guess around and get the answer.

OpenStudy (anonymous):

I mean, have you ever had a teacher tell you, "Guess around till you get the answer"? /:

OpenStudy (anonymous):

Number of dimes: x Number of nickels: (39-x) 10x + 5(39-x) = 270 (cents makes sense) 5x = 75 x = 15, the number of dimes. 24 Nickels checking.... 10*15+5*24 = 270cents or $2.70 THERE YA GO @bearclaw heres the source http://www.algebra.com/algebra/homework/word/coins/Word_Problems_With_Coins.faq.question.879243.html so your final answer is 15 dimes and 24 nickels, bearclaw. Have a nice day! :)